OpenStudy (anonymous):

The Andersons traveled 90 miles in 75 minutes. At this rate, how far would they travel in 2 hours? A. 144 mi B. 165 mi C. 180 mi D. 5400 mi

OpenStudy (welshfella):

work out the speed = 90 / 1.25 mph then use distance = speed * time where speed is as above and time = 2 hours

OpenStudy (welshfella):

i used 1.25 because 75 minutes = 1.25 hours

OpenStudy (welshfella):

ok?

OpenStudy (anonymous):

yes but we are doing 2 hours so then what do you do

OpenStudy (anonymous):

Here u go ;) First convert hours to minute or minute to hours 2 hours= 120 min now apply the formula speed=distance/time rearranging it to solve for distance will be speed times time so: distance= (90/75) x (120 min) distance= 144 miles